For a project, a builder purchased 6 boxes of bolts with each box containing 12 bolts. He ended up finishing the project 9 days early and with 2 bolts left over. How many bolts did he use for the project?
Answer by ikleyn(52790)

.

    6*12 - 2 = 72 - 2 = 70 bolts were used.    ANSWER

Solved, answered and explained. And completed.